OpenStudy (mathmath333):

In the Suniti Building in Mumbai there are 12 floors plus the ground floor.9 people get into the lift of the building on the ground floor.The lift does not stop on the first floor. If 2,3 and 4 people alight from the lift on its upward journey,then in how many ways can they do so?(Assume they alight on different floors)

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & \normalsize \text{ In the Suniti Building in Mumbai there are 12 floors plus }\hspace{.33em}\\~\\ & \normalsize \text{ the ground floor.9 people get into the lift of the building }\hspace{.33em}\\~\\ & \normalsize \text{ on the ground floor.The lift does not stop on the first floor.}\hspace{.33em}\\~\\ & \normalsize \text{ If 2,3 and 4 people alight from the lift on its upward }\hspace{.33em}\\~\\ & \normalsize \text{ journey,then in how many ways can they do so?(Assume}\hspace{.33em}\\~\\ & \normalsize \text{ they alight on different floors)}\hspace{.33em}\\~\\ & a.)\ 11C3 \times 3P3 \hspace{.33em}\\~\\ & b.)\ 11P3 \times 9C4 \times 5C3 \hspace{.33em}\\~\\ & c.)\ 10P3 \times 9C4 \times 5C3 \hspace{.33em}\\~\\ & d.)\ 12C3 \hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

We can select the three floors in 11P3 . Then of the 9 people four can leave in 9C4 ways Three of the remaining 5 can be selected in 5C3 ways so 11p3*9c4*5c3

OpenStudy (mathmath333):

I am having difficulty in interpreting the question

OpenStudy (anonymous):

well you see in 11P3 order is important

OpenStudy (anonymous):

in 9P3 order doesnt matter any way would be fine

OpenStudy (anonymous):

does that make sense ?

OpenStudy (anonymous):

see there is total of 13 floors 12+ ground floor

OpenStudy (mathmath333):

i know order is important in permutation and order isn't important in combination

OpenStudy (mathmath333):

11p3 where this come from ?

OpenStudy (anonymous):

i am coming to that

OpenStudy (anonymous):

because you see they are already on the ground level and they skip the first floor

OpenStudy (anonymous):

so 13 -2 =?

OpenStudy (mathmath333):

11 and 3 ?

OpenStudy (anonymous):

3 is no of ppl leaving the lift

OpenStudy (anonymous):

srry its not 3 ppl its 3 floors

OpenStudy (anonymous):

i was thinking about 5C3

OpenStudy (mathmath333):

ok and how come 9C4

OpenStudy (anonymous):

now selecting 4 ppl from 9 in 9C4

OpenStudy (mathmath333):

why r u selecting only 4 people

OpenStudy (anonymous):

( if 2, 3 and 4 people alight from the lift)*

OpenStudy (anonymous):

now only 8 remaining so now selecting 3 from remaining 5 in 5C3

OpenStudy (anonymous):

ok in 11C3 , 11 floors and the reason i choose 3 floors because (2,3 and 4 ppl alight from the lift)

OpenStudy (dan815):

still need help?

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

about 9C4 and 5C3

OpenStudy (dan815):